[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Feb 10, 2016

Mark DeMaria - NOAA Federal mark.demaria at noaa.gov
Wed Feb 10 12:40:32 UTC 2016


Here is a summary of today's computer operations.

Mark

-----

TAFB:

TSB is continuing to work on getting ftp gate access from the debout
standing desk N-AWIPS system. The required info on the system was provided
to NCO yesterday to open a ticket with OPS-Net to get this done.

NCO is finished configuring the hardware replacement of the derecho N-AWIPS
system at the other standing desk. The new derecho hardware will be swapped
at the standing desk at a time that is convenient for TAFB.

HSU:

No new issues.

Other:

The relocation of the SBN AWIPS cable yesterday at about 2:30 PM for the
roof repair went well. The GOES-West antenna cable was moved at the same
time. There was only minor loss of data during the outage, which lasted
about 1/2 hour.

There was a short internet outage around noon yesterday. TSB will follow up
with NCO on the cause of that.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160210/0311cdbf/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list